# COCOLA: Coherence-Oriented Contrastive Learning of Musical Audio Representations

### v3

@@ V-C MOS vs COCOLA Score
To evaluate COCOLA’s effectiveness, we conducted a subjective test with 41 participants who rated (0-4) the coherence of (pitch) Shifted and (time) Stretched chunks. We also used Random, Ground Truth, Diff-A-Riff and SA ControlNet tracks from the previous experiment. We average the ratings (MOS) for each chunk. Figure 3 shows the correlation plot between MOS and COCOLA Score (using ”COCOLA HPS All”). The Pearson correlation coefficient is $r=$ 0.54 $(p<0.01)$ indicating a good correlation with high statistical signficance.
@@ VI Conclusion
In this paper, we proposed COCOLA, a contrastive encoder for recognizing the coherence between musical stems. Both classification experiments and human scores demonstrate the efficacy of the method. We used COCOLA for benchmarking music accompaniment models, proposing a new evaluation metric for the task. We plan to improve the quality of COCOLA by training on additional stem-level datasets [46] or on pre-separated [13] large scale datasets [43].

### v4

@@ V-C MOS vs COCOLA Score
To evaluate COCOLA’s effectiveness, we conducted a subjective test with 41 participants who rated (0-4) the coherence of (pitch) Shifted and (time) Stretched chunks. We also used Random, Ground Truth, Diff-A-Riff, and SA ControlNet tracks from the previous experiment. We average the ratings (MOS) for each chunk. Figure 3 shows the correlation plot between MOS and COCOLA Score (using ”COCOLA HPS All”). The Pearson correlation coefficient is $r=$ 0.54 $(p<0.01)$, having high statistical significance.
@@ VI Conclusion
In this paper, we proposed COCOLA, a contrastive encoder to evaluate the coherence between musical stems. Classification experiments and human scores demonstrate the effectiveness of the method. We used COCOLA to benchmark music accompaniment models, proposing a new evaluation metric for the task. We aim to improve COCOLA using additional stem-level [50] or pre-separated large-scale [47] datasets.